A conducting loop of face-area A and resistance R is placed perpendicular to a magnetic field B. The loop is withdrawn completely from the field. Find the charge which flows through any cross-section of the wire in the process. Note that it is independent of the shape of the loop as well as the way it is withdrawn.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Area = A, Resistance = R, B = Magnetic field ϕ = BA = Ba cos 0° = BA e = dϕ/dt = BA / 1 ; i = e/R = BA/R ϕ = iT = BA/R